OpenStudy (anonymous):

when n > 5. Prove that for all k. we have

OpenStudy (anonymous):

\[(n-1)! +1\neq n ^{k}\]

myininaya (myininaya):

so we are showing that these two expressions are not equal?

OpenStudy (anonymous):

yea

myininaya (myininaya):

my first thought is find the contrapositive of your statement

OpenStudy (anonymous):

would we prove it by contradiction? we make the equation equal, then prove it somehow. but not sure how to go about it

myininaya (myininaya):

maybe...

OpenStudy (anonymous):

if we sub in n = 6, we would get 121 = 6^k, which doesnt work, but surely its not that simple ?

myininaya (myininaya):

well we want to show it doesn't = for all k

myininaya (myininaya):

let me see if zarkon is here... one sec..

OpenStudy (anonymous):

ok thanks

OpenStudy (across):

Induction?

OpenStudy (anonymous):

not sure induction works its the k which confuses me

myininaya (myininaya):

Show there exist k such that \[(n-1)!+1=n^k\], then \[n \le 5\]. ------------------ So if k=1, then \[(n-1)!+1=n^1=n\] Does (n-1)!+1=n for any n? for n=2 it works: (2-1)!+1=1+1=2=2

OpenStudy (anonymous):

thanks. yea i get what youve done, it works for all n. next step put in n = n+1 ? and not sure how to prove it for all k appreciate the help

myininaya (myininaya):

now let's assume it is true for some integer j so there is some n such that \[(n-1)!+1=n^j\] now let's show the expression is true for some n if j=k+1 so we have \[(n-1)!+1=n^{j+1}\] \[(n-1)!+1=n^j \cdot n \] (n-1)!+1=[(n-1)!+1] \cdot n\] wouldn't that mean n=1 but i'm totally sure how that shows n will be less than 5

myininaya (myininaya):

k=j+1*

OpenStudy (anonymous):

sorry whats the last formula?

myininaya (myininaya):

\[(n-1)!+1=[(n-1)!+1] \cdot n\] 1=n

OpenStudy (anonymous):

thanks

myininaya (myininaya):

but gummage this n is less than 5 but i don't see how this proves what we want to prove we just showed it is true for some n

OpenStudy (anonymous):

we proved that when n < 5 the equation is true, so by contradiction when n > 5 the equation is incorrection, which has proved the question surely? or is there more to it

myininaya (myininaya):

i was trying to prove the contrapositive of your statement if i can do then that means the orginal statement is true i think there is more to it We showed the expression is true for all k for some n we didn't show for which n it is true so the part that is left is to show n will always be less than or =5

myininaya (myininaya):

james what do you think?

OpenStudy (anonymous):

are you sure we're on the right tracks here? lost with this proof now

OpenStudy (jamesj):

I'm thinking about it. (btw, notice it is also false for n = 4.)

OpenStudy (turingtest):

I'm just posting so I will get a notice when this get's solved. I wanna see how it's done :)

OpenStudy (jamesj):

@gummage: what course is this question from?

OpenStudy (turingtest):

I gotta get to school. I hope I can see the solution when I get back. Later.

OpenStudy (anonymous):

sorry back now. still working on it, it's from a Proofs module. It's revision for a test on proofs.

OpenStudy (anonymous):

any progress?

OpenStudy (anonymous):

think ive done it eventually... proving using contrapositive. if you assume k is a natural number and (n−1)!+1= n^k , and sub in n = 6. then you get 121 = 6^k take logs from that log(121) = k*log(6) log(121)/log(6) = k therefore k is not a natural number so it doesn't work. not sure about this but yeah... what you guys think?